### v1

# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

Abstract
Spontaneous reporting system databases are key resources for post-marketing surveillance, providing real-world evidence (RWE) on the adverse events (AEs) of regulated drugs or other medical products. Various statistical methods have been proposed for AE signal detection in these databases, flagging drug-specific AEs with disproportionately high observed counts compared to expected counts under independence. However, signal detection remains challenging for rare AEs or newer drugs, which receive small observed and expected counts and thus suffer from reduced statistical power. Principled information sharing on signal strengths across drugs/AEs is crucial in such cases to enhance signal detection. However, existing methods typically ignore complex between-drug associations on AE signal strengths, limiting their ability to detect signals. We propose novel local-global mixture Dirichlet process (DP) prior-based nonparametric Bayesian models to capture these associations, enabling principled information sharing between drugs while balancing flexibility and shrinkage for each drug, thereby enhancing statistical power. We develop efficient Markov chain Monte Carlo algorithms for implementation and employ a false discovery rate (FDR)-controlled, false negative rate (FNR)-optimized hypothesis testing framework for AE signal detection. Extensive simulations demonstrate our methods’ superior sensitivity—often surpassing existing approaches by a twofold or greater margin—while strictly controlling the FDR. An application to FDA FAERS data on statin drugs further highlights our methods’ effectiveness in real-world AE signal detection. Software implementing our methods is provided as supplementary material.
1 Introduction
Adverse events (AEs) related to drugs and other medical products have a profound impact on public health and have received heightened attention over the past decades. While clinical trials remain the gold standard for assessing medical product safety before approval, they have limitations due to factors such as small sample sizes, ethical considerations, restricted timelines, etc., and the critical need for postmarket monitoring is increasingly recognized. For example, severe AEs such as congenital malformations from Thalidomide and liver injury from the Ketoconazole (Lavertu et al.,, 2021) were identified only after these drugs were approved for the U.S. market. To complement clinical trial findings with real-world evidence (RWE), continuous post-marketing surveillance—a cornerstone of contemporary pharmacovigilance—has proven essential. Accordingly, several Spontaneous Reporting System (SRS) databases have been established globally, including VigiBase (World Health Organization), and FAERS and VAERS (U.S. Food and Drug Administration). These databases collect and manage AE data for regulated medical products, such as drugs, medical devices, vaccines, and therapeutic biologics. For exposition, we focus only on drugs hereinafter, though analogous considerations apply to other medical products monitored in SRS databases.
The SRS databases serve as key resources for assessing real-world evidence in postmarket drug safety by cataloging a large number of case safety reports linking drugs to relevant AEs. For example, the FAERS database has compiled over $29$ million reports by 2024. These databases are inherently observational and suffer from selection bias, measurement error, confounding, and missing data, among other factors (Markatou and Ball,, 2014), precluding causal inference. Instead, the focus is on detecting signals: identifying drug-AE pairs that appear disproportionately more often than their null expected counts, derived under the assumption of independence between the drug and the AE using their marginal totals and the dataset’s grand total counts. Association analyses then estimate and assess the significance of signal strengths, quantifying the differences between the observed and (null) expected counts or related quantities defined by parameters of some underlying statistical model.
Numerous SRS data mining and signal detection approaches have been developed over the past decades. Examples include the proportional reporting ratio (PRR) (Evans et al.,, 2001), the reporting odds ratio (ROR) (Rothman et al.,, 2004), formal frequentist methods such as the likelihood ratio test (LRT) and its variants (Huang et al., 2011; Chakraborty et al., 2022; Ding et al., 2020; and the references cited therein), and Bayesian methods such as the Gamma Poisson Shrinker (GPS) (DuMouchel,, 1999), Bayesian confidence propagation neural network (BCPNN) (Bate et al.,, 1998), and Dirichlet process (DP) Poisson Gamma mixture model (Hu et al.,, 2015). Some approaches—e.g., PRR and ROR—employ ad-hoc thresholds on empirical signal strength estimates, while the more formal methods—e.g., LRT and Bayesian approaches—employ rigorous probabilistic tools for statistically guaranteed signal detection, including controlled type I errors and false discovery rates (FDR). Despite these advances, signal detection for rare AEs and newer drugs remains challenging due to their small (marginal total) sample sizes, which leads to small observed and expected report counts in individual drugs, reducing the statistical power to detect the underlying signals. Yet, a large and growing body of research has documented the critical need for reliable pharmacovigilance signal detection precisely under these setups (see, e.g., Poluzzi et al., 2012; Sardella and Belcher, 2018; Raschi et al., 2021, and the references cited therein).
Bayesian methods facilitate principled inference and signal detection for SRS data by modeling drug-AE signal strengths as model parameters with suitable priors and utilizing their posterior distributions. However, prior selection is crucial. An overly restrictive prior—e.g., a single gamma shrinkage prior for all drug-AE signal strength parameters—may cause excessive shrinkage, pulling estimates toward a single common value (Hu et al.,, 2015). Conversely, an overly flexible prior—e.g., independent flat priors for all parameters—may inadequately regularize their estimation, resulting in overfitting and noisy estimates. State-of-the-art Bayesian data mining approaches balance flexibility and regularization using semi/nonparametric shrinkage priors—e.g., a two-component gamma mixture prior (the GPS model of DuMouchel, (1999)) or a DP mixture prior with a base gamma distribution (Hu et al.,, 2015)—for AE signal strengths separately for each drug. This stabilizes signal detection, provided the drugs are treated either all independently or as a single group.
However, real-world SRS databases record drugs with nuanced similarities in their AEs. For instance, various statin drugs, used to lower blood cholesterol and cardiovascular risk, share chemical compositions and function, leading to common AEs like rhabdomyolysis (Montastruc,, 2023). On the other hand, small differences in drug composition and function can cause distinct AEs or variations in AE severity; e.g., myoglobinuria is typically associated with Atorvastatin but not Lovastatin (Seneff,, 2016). Figure 1 shows the pairwise Spearman correlation coefficients for occurrences of 46 common AEs of six statin drugs as cataloged in FAERS (Q3 2014-Q4 2024). These correlations reveal strong yet nuanced between-drug associations: they are consistently above 0.70, ruling out independence, yet many are noticeably below one, suggesting against a single grouping of all drugs.
Figure 1: Spearman correlation matrix for occurrence counts of 46 common AEs (PTs) among statin drugs in the FDA FAERS data (2014 Q1-2020 Q4)
Existing SRS data mining approaches, both Bayesian and frequentist, rarely account for nuanced between-drug associations. Instead, they often assume either independence of drugs (the original LRT, BCPNN, GPS, and DP Poisson-Gamma) or a single group for all drugs (extended LRT of Huang et al., (2011)). These restrictive assumptions limit the statistical power and sensitivity for detecting signals, particularly for rare AEs and newer drugs with small observed and expected counts. Indeed, our extensive simulations (Section 4) show that existing methods suffer from poor sensitivity (after controlling FDR) when detecting signals from datasets containing multiple but weak true signals, rare AEs, and newer drugs.
We address this critical gap with a novel local-global mixture DP-based Bayesian method that balances flexibility and shrinkage in AE signal strengths for each drug while enabling data-adaptive information sharing across drugs based on their associations, enhancing statistical power for signal discovery. In this framework, a local DP models distinct AE signal strength patterns specific to each drug while a global process defines a common pattern across drugs; the final AE signal strengths for a drug are obtained from a probabilistic mixture of these local and global processes. By appropriately controlling the local-global contributions, our approach can flexibly capture complex between-drug associations in SRS data.
For full Bayesian inference from our model, we employ weakly informative priors customized for SRS data mining on various DP hyperparameters. This eliminates the need for any additional tuning and acknowledges all modeling uncertainty, including hyperparameters. Subsequently, we develop efficient Markov chain Monte Carlo (MCMC) algorithms for posterior computation. Signal detection is performed using a rigorous hypothesis testing framework that controls the false discovery rate (FDR) while minimizing the false negative rate (FNR), both estimated via MCMC-computed posterior probabilities. Our default implementation supports a wide range of data-generating settings for robust signal detection. Additional inferential goals, such as identifying structural zero (physically or biologically impossible drug-AE pairs), can be seamlessly incorporated through additional model layers without compromising signal detection accuracy, which we also discuss.
Our key contributions are as follows. First, we propose local-global mixture DP models for SRS signal detection (Section 3). Local-global priors are widely used in contemporary statistics (Bhadra et al., 2019 and references cited therein) and have been employed in DP modeling of grouped data, notably via hierarchical DP models (Teh et al.,, 2004). Our framework instead adopts a more transparent two-component mixture of DPs (Antoniak,, 1974) for local-global pattern decomposition. To our knowledge, the use of local-global DP models for principled SRS data mining is novel. We develop efficient MCMC algorithms for model implementation (Section 3.3 and Supplement S.3; software in the Supplement) and leverage FDR-controlled, FNR-optimized hypothesis tests for signal detection (Section 3.4). Second, we provide an intuitive overview of DP-based SRS data mining models to aid broader understanding (Section 3.1). Third, through extensive simulations (Section 4), we show that our methods achieve comparable or superior sensitivity to existing approaches (reviewed in Supplement S.1) when only a few strong signals are present across common AEs but yield substantial sensitivity gains—often twofold or more—when many weak signals exist among rare AEs and newer drugs, all while maintaining strong FDR and type I error control. Finally, our analysis of a real-world statin dataset (Section 5) from the FDA FAERS database demonstrates our methods’ ability to effectively detect AE signals in practice.
2 SRS DATA, OUR NOTATION, AND PROBABILITY MODELS
An SRS database stores a large number of individual case safety reports that can be summarized as a high-dimensional contingency table cataloging the occurrences of multiple AEs across the rows and various drugs across the columns. Let $I$ and $J$ denote the total number of AEs and drugs under consideration, respectively. Further, let $n_{ij}$ be the total number/count of reported cases for drug-$j$ and AE-$i$; $i=1,\dots,I$ and $j=1,\dots,J$ constituting the contingency table $((n_{ij}))$. Define $n_{i\bullet}=\sum_{j=1}^{J}n_{ij}$, $n_{\bullet j}=\sum_{i=1}^{I}n_{ij}$, and $n_{\bullet\bullet}=\sum_{i=1}^{I}\sum_{j=1}^{J}n_{ij}$.
Signal detection in SRS contingency tables often relies on the observed relative reporting rates, $n_{ij}/E_{ij}$, or a related quantity, as a measure of signal strength. Here, $E_{ij}=n_{i\bullet}n_{\bullet j}/n_{\bullet\bullet}$ represents the null expected count of drug-$j$ and AE-$i$, assuming independence between the drug and AE. Intuitively, if $n_{ij}/E_{ij}\gg 1$, the observed frequency for $(i,j)$ is much higher than the null expected count, suggesting a potential signal. Challenges arise for small $n_{\bullet j}$ (e.g., a new drug) and/or small $n_{i\bullet}$ (rare AE), leading to a small $E_{ij}$ and rendering $n_{ij}/E_{ij}$ to be noisy. Bayesian approaches parameterizing $n_{ij}/E_{ij}$ (or an analogous quantity) using a probability model for $n_{ij}$ can stabilize inference on signal strengths under these setups by incorporating suitable prior distributions that can facilitate principled information sharing across AEs and drugs, leading to improved signal detection.
The Poisson model $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$ is commonly used for its analytical simplicity and computational tractability. Here the signal strength parameter $\lambda_{ij}$ parametrizes $n_{ij}/E_{ij}$, with the pair $(i,j)$ considered a signal if $\lambda_{ij}\gg 1$ and non-signal otherwise. When the Poisson assumption is violated, particularly due to an excess of zero counts arising from structural zeros—i.e., drug-AE pairs that are physically or biologically impossible to co-occur—a zero-inflated Poisson (ZIP) model is often preferred. In this framework, $n_{ij}$ is modeled as a mixture of two components: a $\text{Poisson}(\lambda_{ij}E_{ij})$ component for drug-AE pairs that can co-occur and a point mass at 00, denoted by $\delta_{\{0\}}$ (the Kronecker delta), for structural zeros. Importantly, the structural zero positions are typically unknown and inferred from the SRS dataset.
A Bayesian SRS data mining approach begins with such a probability model for $n_{ij}$ (e.g., Poisson or ZIP), yielding a likelihood for the signal strengths $\lambda_{ij}$ that is combined with a prior distribution to derive the posterior distribution of $\lambda_{ij}$ for inference. The marginal distribution of $n_{ij}$, after integrating over the prior, can be more flexible than the likelihood model for $n_{ij}$. Indeed, with an appropriate nonparametric mixture prior, a flexible marginal distribution accommodating ZIP-like behavior can be achieved using only a Poisson likelihood for $n_{ij}$ (see Remark Remark under Section 3). Thus, Bayesian SRS data mining methods are less sensitive to the likelihood choice (Poisson or ZIP) when a flexible prior is used. However, if the zero-inflation probability parameters are of interest, the ZIP model remains useful.
3 NOVEL LOCAL-GLOBAL DP MODELS FOR SRS DATA MINING
We begin in Section 3.1 with a review of the drug-specific DP Poisson model of Hu et al., (2015), outlining its mechanism and providing guidance on the hyperparameters. We then introduce our novel local-global DP framework (Section 3.2) that accommodates between-drug associations and enables effective information sharing across drugs.
3.1 A conceptual overview on the DP Poisson model
Consider the Poisson likelihood $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$, with the DP prior (Hu et al.,, 2015):
$\lambda_{ij}\sim D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}};\ D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ (1)
for the signal strength parameters $\{\lambda_{1j}\dots,\lambda_{Ij}\}$ associated with each drug $j$. Here $\delta_{\{\theta\}}$ denotes a degenerate distribution at $\theta$ and DP$(\alpha,G(\beta))$ is a Dirichlet process with concentration hyperparameter $\alpha>0$ and base distribution $G(\beta)\equiv$ Gamma(shape = $\beta$, rate = $\beta$) with hyperparameter $\beta>0$. The discrete mixture structure ensures that $D_{j}$, the prior for $\lambda_{ij}$, is almost surely discrete/categorical placing probability weights $\eta_{hj}$; $10\leq\eta_{hj}\leq 1;\sum_{h=1}^{\infty}\eta_{hj}=$ 1 on atoms $\theta_{hj}$. This permits clustering, allowing multiple $\lambda_{ij}$ to share an atom $\theta_{hj}$ as a common value. The probability weights $\eta_{hj}$ are generated using stick-breaking (Sethuraman,, 1994):
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj})\text{ for }h\geq 2;\ v_{hj}\sim\text{independent Beta}(1,\alpha_{j})\text{ for }h\geq 1$ (2)
based on the DP concentration hyperparameter $\alpha_{j}>0$. Meanwhile, the atoms are generated independently from the DP base distribution: $\theta_{hj}\sim G(\beta_{j})$. In practice, the infinite mixture defining $D_{j}$ can be well-approximated by a finite mixture $\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$ with a sufficiently large $K$ (Ishwaran and Zarepour,, 2002). This approximation simplifies model fitting and offers intuitive insights on the hyperparameters $\alpha_{j}$ and $\beta_{j}$, guiding their specification.
Role of $\alpha_{j}$. The stick-breaking representation (2) elucidates that when $\alpha_{j}<1$, each $v_{hj}$ is, on average, greater than $1/2$ with smaller $\alpha_{j}$ favoring $v_{hj}$ closer to one. Consequently, $\eta_{hj}$ decreases sharply in $h$, leading the mixture (1), or its finite approximation with $K$ components, to retain only a few non-empty components with sufficiently large weights $\eta_{hj}$. Under a finite mixture approximation with $K_{+}<K$ non-empty components, (1) models $\lambda_{ij}$ using a categorical distribution with $K_{+}$ categories, placing point masses at the atoms $\theta_{1j},\dots,\theta_{K_{+}j}$ with weights $\eta_{1j},\dots,\eta_{K_{+}j}$. This fosters clustering, restricting $\lambda_{1j},\dots,\lambda_{Ij}$ to at most $K_{+}$ distinct values. When $K_{+}<I$, multiple $\lambda_{ij}$ share the same cluster $h$ admitting the common value $\theta_{hj}$, facilitating their pooling for information sharing. While a smaller $\alpha$ encourages a smaller $K_{+}$ aiding greater information sharing, a larger $K_{+}$ ($>2$) is permitted in the posterior when supported by data. This contrasts with the GPS framework of DuMouchel, (1999) commonly used for SRS signal detection (see Section S.1 for a review) which strictly categorizes $\lambda_{ij}$ into two groups, intuitively, signal ($\lambda_{ij}\leq 1$) and non-signal ($\lambda_{ij}\gg 1$). The DP framework, by allowing $K_{+}>2$, enhances heterogeneity acknowledgment within both signal and non-signal groups. Figure 2a illustrates this with $I=6$ AEs and $K_{+}=3$ clusters.
Because a smaller $\alpha_{j}$ promotes greater information sharing, a prior assigning sufficient mass to small positive values is recommended for stable estimation of $\lambda_{ij}$. Hu et al., (2015) used a Uniform(0.2, 10) prior for $\alpha_{j}$. We suggest a more aggressive weakly informative default prior, viz., Exponential(rate = $\psi_{\alpha}$) with a moderately large $\psi_{\alpha}$, e.g., $\psi_{\alpha}=3$ (placing $\approx 95\%$ mass in $(0,1)$), which favors small $\alpha_{j}$ while allowing larger values if needed. This choice also ensures conjugacy for efficient Gibbs sampling updates for MCMC implementation.
Figure 2: Schematic diagrams visualizing the mechanisms of a “local only” DP prior with one drug (Panel A) and local-global mixture DP prior with $J=3$ drugs (panel B). Throughout, darker shades represent larger values. Panel A: The observed relative reporting rates $n_{ij}/E_{ij}$ for $I=6$ AEs for the drug $j$ (Panel A.1) show high variability across $i$ (different shades of green) due to randomness in observed data. The atoms $\theta_{hj}$ (Panel A.2) of the DP prior enable clustering of $I=6$ AE signal strengths into $K_{+}=3$ non-empty DP clusters. This ensures information sharing: the AEs $(i=1,i=3,i=5)$ in cluster 1 and separately the AEs $(i=2,i=6)$ in cluster 2 (cluster memberships are visualized via arrows between Panels A.1 and A.2) share their observed $(n_{ij},E_{ij})$ information to inform their cluster-specific DP atoms (single atom per cluster). These DP atoms, together with the cluster memberships, produce the final $I=6$ AE signal strengths $\lambda_{ij}$ (Panel A.3). Panel B: The noisy observed data $(n_{ij},E_{ij})$ leading to noisy observed relative reporting rates $n_{ij}/E_{ij}$ (Panel B.1; different shades of green) are used to produce “local” signal strengths $\lambda_{ij}^{l}$ for $I=6$ AEs and $J=3$ drugs using a separate, independent local DP prior for each drug (Panel B.2.1), and global AE signal strengths $\lambda_{i}^{g}$ for shared by all $J=3$ drugs based on a common global DP prior (Column B.2.2). Probabilistic two-component mixtures of $\lambda_{ij}^{l}$ and $\lambda_{i}$ (Panel B.3) based on the local distribution indicators $z_{ij}^{l}$ produce the final realized signal strengths $\lambda_{ij}$ (Panel B.4).
Role of $\beta_{j}$. The parameter $\beta_{j}$ controls variability among the atoms $\theta_{hj}$ drawn from the base $G(\beta_{j})\equiv$ Gamma(shape = $\beta_{j}$, rate = $\beta_{j}$) distribution, thereby controlling the shrinkage in $\lambda_{ij}$. Since $G(\beta_{j})$ has unit mean and variance $1/\beta_{j}$, larger (smaller) values of $\beta_{j}$ encourage all $\theta_{hj}$ to concentrate more (less) around $1$, inducing more (less) shrinkage in $\lambda_{ij}$ (around $1$).
[…]
Remark.
[…]
3.2 The local-global DP Poisson model
The DP model in Section 3.1 enables flexible and effective signal detection for each drug separately but does not account for associations between drugs in their signal strenghts. To capture these associations, we introduce the concept of local and global (prior) distributions for AE signal strengths. Specifically, we assume that the realized AE signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for each drug $j$ is drawn from a mixture of a “local” distribution specific to the drug and a “global” distribution shared among drugs with similar compositions or functions.
Under the same likelihood $n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$, we consider the following local-global mixture of DP prior for $\lambda_{ij}$:
$\lambda_{ij}=z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g};\ \lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D;\ z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
$D_{j}\sim DP(\alpha_{j},G(\beta_{j}));\ D\sim DP(\alpha,G(\beta)).$ (3)
Here, the realized signal strengths $\{\lambda_{1j},\dots,\lambda_{Ij}\}$ for a drug $j$ is generated by combining local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$ drawn from a drug-$j$-specific local distribution $D_{j}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$ as in (1) and global strengths $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ generated from a global distribution $D\sim\text{DP}(\alpha,G(\beta))$ shared by all drugs. The binary local distribution indicators $z_{ij}^{l}$ governed by the common local proportion $\pi^{l}$, determine the probabilistic mixing of $\lambda_{ij}^{l}$ and $\lambda_{i}^{g}$.
We note that the proposed local-global mixture DP prior (3) enables information sharing in signal strengths across both AEs and drugs. Specifically, for each drug $j$, the realized signal strengths $\{\lambda_{i^{\prime}j}:\text{all }i^{\prime}\text{ with }z_{i^{\prime}j}=1\}$ exhibit the same $D_{j}$-induced clustering (see Section 3.1) as the local strengths $\{\lambda_{1j}^{l},\dots,\lambda_{Ij}^{l}\}$, facilitating information sharing across AEs. Meanwhile, for any AE $i$, the signal strengths $\{\lambda_{ij^{\prime}}:\text{all }j^{\prime}\text{ with }z_{ij^{\prime}}=0\}$ form a cluster attaining the common global value $\lambda_{i}^{g}$, and enable information sharing across drugs (the global values $\{\lambda_{1}^{g},\dots,\lambda_{I}^{g}\}$ themselves admitting clustering via $D$). Figure 2B illustrates this mechanism using $I=6$ AEs and $J=3$ drugs. This bi-level clustering contrasts with the single-level (AE-only) clustering induced by the DP prior of Hu et al., (2015). As demonstrated in extensive simulations (Section 4), the bi-level clustering and information sharing play a critical role in stabilizing the posterior and enhancing signal detection for $\lambda_{ij}$, particularly for drug-AE pairs where $n_{i\bullet}$ is small (rare AEs) and/or $n_{\bullet j}$ is small (newer drugs), leading to small $E_{ij}$.
The drugs under consideration should be similar in composition or function to justify a single global distribution $D$. If distinct drug groups exist with no similarities, separate global processes can be used to enable separate within-group information sharing. Additionally, SRS data often include a collapsed, heterogeneous comparator category $J$ (e.g., “Other drugs”), serving as a baseline group with few, if any, distinct AE signals. In such cases, shrinking all $\lambda_{iJ}$ toward $1$ using a prior such as $\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$ with an appropriate hyperparameter $\tau$ can improve signal detection and inference (see Supplement S.2.1).
As noted in Remark Remark, a separate ZIP likelihood $n_{ij}\sim\text{ZIP}(\lambda_{ij}E_{ij},\omega_{j})$ is unnecessary for signal detection due to the flexibility of the DP prior. However, it may be used if direct inference on structural zero positions/probabilities is required. Complete hierarchical Bayesian representations of our Poisson and ZIP-based models are provided in Supplement S.2.1.
Remark.
The Hu et al., (2015) model can be viewed as a “local-only” special case of our local-global model, with only drug-specific local distributions $D_{j}$ for $\lambda_{ij}$ and no global distribution (i.e., $\pi^{l}=1$). At the other extreme, a “global-only” model ($\pi^{l}=0$) yields common $\lambda_{i}$ across drugs. This is a restrictive model that cannot differentiate between drugs and only permits “global” signal detection—assessing whether an AE-$i$ is a signal for any drug under consideration (see Supplement S.2.2 for further discussion on this model).
Remark.
Extensive literature exists on local-global DP models for sharing information across groups while preserving DP-based clustering within each group. A common approach employs hierarchical DPs (Teh et al.,, 2004), where local DPs share a global DP as a base distribution. In contrast, our approach leverages a two-component DP mixture framework (Antoniak,, 1974), separately modeling local and global AE signal strengths and thus offering a more transparent bi-level clustering over AEs and drugs, enabling more flexible estimation of the local and global parameters for greater flexibility in SRS applications.
Remark.
In this paper, we focus on drug-based pharmacovigilance, identifying AE (row) signals given one or more drugs (columns in SRS contingency tables). An alternative AE-based approach discovers drugs strongly associated with a given set of AEs (Ding et al.,, 2020; Liu et al.,, 2024). A row-based version of our column-based prior (3), obtained by switching the roles of $i$ and $j$, can be used for such analyses.
3.3 Full Bayesian implementation using MCMC
For full Bayesian inference on our models, we place independent weakly informative proper priors on all model hyperparameters. Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 3.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform$(0,1)$ priors for the probability weight parameters, viz., $\pi_{l}$ and $\omega_{j}$ (if included in the model).
We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components. Each MCMC iteration involves data augmentation to generate latent DP mixture allocations, followed by Gibbs sampling updates for the stick-breaking weights and atoms and the concentration hyperparameters for each local and global DP. The local distribution indicators $z_{ij}^{l}$, local distribution proportion $\pi^{l}$, AE signal strengths $\lambda_{iJ}$ for the reference drug $J$, and zero-inflation indicators $y_{ij}$ and probabilities $\omega_{j}$ (if included in the model) also admit standard Gibbs updates. However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ stepping-out slice sampling (Neal,, 2003). Detailed MCMC steps are provided in Supplementary Algorithms S.1, S.2 and S.3. For signal detection, we use an FDR-controlled, FNR-optimized hypothesis testing framework based on posterior draws for $\lambda_{ij}$, described next.
3.4 Signal detection using hypothesis testing
For signal detection we consider testing the hypotheses:
$H_{0,ij}:\lambda_{ij}\leq 1\text{ vs. }H_{1,ij}:\lambda_{ij}>1$
for all AEs $i=1,\dots,I$ and all drugs $j=1,\dots,J$. The null hypothesis $H_{0,ij}$ is rejected in favor of $H_{1,ij}$, identifying the drug-$j$ and AE-$i$ pair to be a signal, if $T_{ij}>k$, where $T_{ij}$ is a test statistic derived from the posterior distribution and $k>1$ is an appropriate threshold. A common choice for $T_{ij}$ is the marginal posterior posterior quantile $T_{ij}(p)$, defined by $\Pr{\!}_{\text{posterior}}(\lambda_{ij}\leq T_{ij}(p))=p$ for a small pre-specified quantile level $p\in(0,1)$, where $\Pr{\!}_{\text{posterior}}(\cdot)$ denotes posterior probability given the data, and can be computed via MCMC draws. Hu et al., (2015) suggest using the 5-th posterior quantiles $T_{ij}(0.05)$ as the test statistic with a rejection threshold of $k=2$; however, such deterministic thresholds can be overly conservative (curtailing sensitivity) or overly liberal (failing to control false discoveries). Instead, we consider a family of tests $\{T_{ij}(p)>k\}$ with varying $p$ (reasonably small) and $k>1$, selecting the optimal $(\hat{p},\hat{k})$ that controls the false discovery rate (FDR) while minimizing the false negative rate (FNR). Given $p$ and $k$, FDR and FNR are estimated using the posterior distribution as (Ahmed et al., 2010; Müller et al., 2006; in the expressions below $1(A)$ denotes the indicator of a set $A$, $\hat{=}$ denotes “estimated by”, and $T^{\text{obs}}$ is the computed value of $T$):
$\text{FDR}(p,k)\ \hat{=}\ \frac{\sum_{i,j}q_{ij}\ I(T_{ij}(p)^{\text{obs}}>k)}{\sum_{ij}1(T_{ij}(p)^{\text{obs}}>k)},\quad\text{FNR}(p,k)\ \hat{=}\ \frac{\sum_{i,j}(1-q_{ij})\ 1(T_{ij}(p)^{\text{obs}}\leq k)}{\sum_{ij}1(T_{ij}(p)^{\text{obs}}\leq k)},$ (4)
where $q_{ij}=\Pr{\!}_{\text{posterior}}(\lambda_{ij}\leq 1)$ is the posterior probability of $H_{0,ij}$, computable from MCMC draws. A grid search is used to determine the optimal $(\hat{p},\hat{k})$ that ensures $\text{FDR}(\hat{p},\hat{k})\leq\alpha$ for a pre-specified level $\alpha$ while minimizing FNR (see Supplement S.3.3 for details on the grids).
In practice, drug-AE pairs with $n_{ij}=1$ are often deemed non-signals, even if $\lambda_{ij}\gg 1$ with high posterior probability, leading to $T_{ij}(p)>k$ for suitably chosen $p$ and $k$. To prevent their detection as signals, the rejection region can be modified to $\{T_{ij}(p)>k\text{ and }n_{ij}>1\}$.
4 SIMULATION ASSESSMENTS
In this section, we conduct extensive simulations to assess the signal detection performance of our local-global DP Poisson and ZIP models under a wide array of data-generating settings. To generate random SRS contingency tables of dimension $I\times J$, we first define a prespecified true signal strength matrix (described below) $\Lambda^{0}=((\lambda_{ij}^{0}))$ encoding signals, structural zeros, non-signals, and between-drug dependency patterns. Using a reference $I\times J$ contingency table (described next) from a real-world SRS database, we extract row and column totals $\{\bm{n}_{i\bullet}:i=1,\dots,I\}$ and $\{\bm{n}_{\bullet j}:j=1,\dots,J\}$, respectively, to facilitate random contingency table generation. Next, we generate $\bm{p}_{i\bullet}\sim\text{Dirichlet}(\bm{n}_{i\bullet})$ and $\bm{p}_{\bullet j}\sim\text{Dirichlet}(\bm{n}_{\bullet j})$ for all $i$, $j$ defining the random probability matrix $\bm{p}=((p_{ij}))$ with $p_{ij}\propto\lambda_{ij}^{0}\bm{p}_{i\bullet}\bm{p}_{\bullet j}$ and $\sum_{i,j}p_{ij}=1$. Finally, we generate a random SRS contingency table by sampling from $\text{Multinomial}(\bm{n}_{\bullet\bullet};\bm{p})$, collapsing any generated $n_{ij}=1$ corresponding to a true signal (as defined by $\lambda^{0}_{ij}$) to $n_{ij}=0$ afterward to eliminate signal detection at $n_{ij}=1$.
Reference SRS dataset to compute row and column totals. We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,610 reported cases (Chakraborty et al.,, 2023), to construct our reference SRS contingency table. We focus on six statin medications: Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, collapsing all other drugs into a comparator “Other drugs” category. Using 46 preferred terms (PTs) for AEs, we form a $47\times 7$ contingency table summarizing the report counts for each drug-AE pair. This dataset has been previously analyzed in Chakraborty et al., (2022).
Constructing the true signal strength matrix $\Lambda^{0}$. To generate $\Lambda^{0}$, we define two signal patterns: (1) fixed row signals, where all drugs (columns) exhibit signals for the same AEs (PTs; rows), and (2) random signals, where signals appear on randomly chosen AEs (different row for each drug). A larger number of fixed row signals increase the between-drug associations, measured via correlations between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ for different statin drugs $j$ and $j^{\prime}$, while more random signals weaken them. The matrix $\Lambda^{0}$ is constructed by specifying fixed row signals and embedding additional random signals in the remaining rows across all statin columns. Signal strengths in these cells range from $1<\lambda_{ij}^{0}\leq 3$ (see Table 1). The remaining cells $(i^{\prime},j^{\prime})$ are designated as non-signals with $\lambda_{i^{\prime}j^{\prime}}^{0}=1$. To introduce zero inflation, we prespecify true $\omega_{j}^{0}$ (00 or $0.10$; Table 1), and set $\omega_{j}^{0}I$ randomly selected non-signal cells to $\lambda_{ij}^{0}=0$. This approach yields a controlled $\Lambda^{0}$ to generate random SRS contingency tables. Within this setup, we designed three simulation scenarios, each with and without zero-inflation (see Table 1).
Table 1: Simulation scenarios describing the construction of the true signal strength matrix $\Lambda^{0}$ encoding signals, non-signals, zero-inflation, and between-drug associations. The first two rows in the table show the number of fixed-AE/row signals and the number of random signals per (statin) column in $\Lambda^{0}$. The zero inflation rate (third row in the table) shows the drug-specific true zero-inflation probability $\omega_{j}^{0}$ in each simulation scenario. The comparator “Other drugs” column has all non-signal $(\lambda_{ij}^{0}=1)$ cells. Finally, the fourth row in the table shows the average Pearson’s correlation coefficient $\rho$ between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ over all statin drug pairs $j$ and $j^{\prime}$ (excluding the “Other drugs” category) as a measure of pairwise dependence between drugs.
Simulation Scenario 1a 1b 2a 2b 3a 3b
No. of fixed row signals 1 1 3 3 6 6
No. of random signals per column 5 5 3 3 1 1
Zero-inflation rate per column 0 0.1 0 0.1 0 0.1
Average $\rho$ between columns 0.04 0.04 0.43 0.43 0.83 0.83
  • The common/fixed PTs (fixed AEs/rows) that are considered signal in each case are as follows:
    • Case 1: Rhabdomyolysis
    • Case 2: Muscle Disorder, Myalgia, Rhabdomyolysis
    • Case 3: Blood Creatine Phosphokinase Increased, Muscle Disorder, Muscle Rupture, Musculoskeletal Discomfort, Myalgia, Rhabdomyolysis
  • A common signal strength value $\lambda^{0,\text{signal}}$ is assigned to all signal cells in a given simulation setting, and we vary $\lambda^{0,\text{signal}}\in\{1.1,1.2,1.3,1.4,1.5,1.7,1.9,2,2.5,3\}$.
Replicated Data Generation, Model Fitting, and signal detection. For each constructed signal strength matrix $\Lambda$ (Table 1), and using the random contingency table generation mechanism described earlier, we generate $R=1{,}000$ independent random contingency table replicates. In each replicate, we fit our models via MCMC sampling (Supplementary Algorithms S.1, S.2 and S.3) with $10{,}000$ post-burn-in draws after discarding the first $5,000$ burn-in draws; see Supplement S.3 for additional computational details, including starting points. From the computed posteriors, we conduct signal detection using hypothesis testing (Section 3.4).
We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al., (2015) (henceforth DP Hu et al). For our models and DP Hu et al., signal detection uses an optimized posterior quantile level $\hat{p}$ (fixed at 0.05 for DP Hu et al.) and threshold $\hat{k}$ ensuring FDR $\leq 0.05$ while minimizing FNR (Section 3.4), with the additional constraint $n_{ij}>1$ for our methods. To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted (Müller et al.,, 2006) posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$. Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT (Chakraborty et al.,, 2022; Huang et al.,, 2011; Ding et al.,, 2020), which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present (see Supplement S.1 for a detailed review of competing approaches).
We assess signal detection for all methods using replication-based FDR, sensitivity, average Type I error, and F-score (Figure 3). For each method, these metrics are computed as FDR = $R^{-1}\sum_{r}\bigl{(}\text{FP}_{r}/(\text{FP}_{r}+\text{TP}_{r})\bigr{)}$, sensitivity = $R^{-1}\sum_{r}\bigl{(}\text{TP}_{r}/(\text{TP}_{r}+\text{FN}_{r})\bigr{)}$, average type I error = $R^{-1}\sum_{r}\bigl{(}\text{FP}_{r}/\text{TN}_{r}\bigr{)}$, and F-score = $R^{-1}\sum_{r}\bigl{(}2\text{TP}_{r}/(2\text{TP}_{r}+\text{FN}_{r}+\text{FP}_{r})\bigr{)}$ where $\text{TP}_{r}$ and $\text{FN}_{r}$ denote the number of true signals that are discovered and not discovered respectively, and $\text{FP}_{r}$ and $\text{TN}_{r}$ are the number of true non-signals discovered and not discovered in replicate $r$ based on the signal strength values as cataloged in the corresponding true $\Lambda$.
Figure 3: Simulation results: the horizontal axis shows the true signal strength $\lambda_{0}$ for signal cells, the vertical axis shows the value of evaluation metrics. Each column presents one metric and each row presents one simulation scenario in Table 1.
The first two rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the between-drug association is near zero (Cases 1a,b). Under moderate between-drug association ($\rho=0.43$; Cases 2a,b), our methods perform noticeably better in detecting more true signals, achieving up to twice the sensitivity, particularly for weak or moderate true signal strengths ($1.1\leq\lambda^{0,\text{signal}}\leq 1.7$). The gain is even further pronounced under high between-drug association ($\rho=0.83$; Cases 3a,b), where our methods attain sensitivities of $\geq 0.70$ with only moderate true signal strength ($\lambda^{0,\text{signal}}\geq 1.7$), while all competitors fail to reach $0.70$ sensitivity even for $\lambda^{0,\text{signal}}=3$. These results highlight the substantial boost in signal detection from effectively leveraging between-drug similarities as facilitated by our models. In all settings, the FDR and average type I error is strictly controlled for all methods.
The Poisson and ZIP-based DP local-global models perform similarly across all scenarios regardless of the presence of zero-inflated observed counts (sub-cases a and b), due to the flexibility of the DP-based nonparametric framework (see Remark Remark).
All computations were performed on the cluster computing facility at the Center for Computational Research at the University at Buffalo, with each process running on 16 GB RAM and a clock speed of 1.6-3 GHz. The median computation time per replication for our DP local-global methods (both Poisson and ZIP) was 145 seconds.
5 REAL DATA ANALYSIS
We use the statin data from Section 4 to demonstrate our method’s utility in real-world SRS data mining. Applying our DP local-global models (using the same MCMC and test statistic generation settings as used for simulated data analysis in Section 4) we compare detected signals against those from the DP Hu et al., BCPNN, GPS, and LRT. Table 2 presents the number of signals identified across statin drugs. The results indicate identical performance for our Poisson and ZIP models, detecting more signals than all competing methods.
Table 2: Summary table of discoveries from our proposed DP local-global models, and comparing them to the other existing methods. The count in each cell represents the number of AE signals detected by the method (along the columns), first separately for each drug and then in total (rows). The number of discoveries made uniquely by our DP Poisson local-global method but not by the other methods are presented in parentheses.
Drug DP Poisson DP ZIP DP Hu et al. BCPNN GPS LRT
Atorvastatin 33 33 (0) 33 (0) 28 (5) 26 (7) 28 (5)
Fluvastatin 14 14 (0) 13 (1) 10 (4) 12 (2) 12 (2)
Lovastatin 12 12 (0) 10 (2) 9 (3) 9 (3) 9 (3)
Pravastatin 23 23 (0) 17 (6) 11 (12) 11 (12) 11 (12)
Rosuvastatin 29 29 (0) 28 (1) 25 (4) 25 (4) 25 (4)
Simvastatin 29 29 (0) 28 (1) 26 (3) 25 (4) 25 (4)
Total 140 140 (0) 129 (11) 110 (31) 110 (32) 110 (30)
To further examine the results, we focus on 20 PTs (AEs/$i$) with the highest average (across statin drugs) posterior medians of $\lambda_{ij}$ from our DP Poisson local-global model (Figure 4). The heatmap indicates non-detection (gray) or the degree of information sharing between drugs, measured via the posterior probabilities of $1-z_{ij}^{l}$ (white-blue gradient; darker blue signifies greater information sharing at each AE). Our method detects several rare AEs with $n_{ij}=2$, some also identified by certain competing methods (though no other method detects all these AEs simultaneously) and some not detected by any competing method (not shown in Figure 4; see Table 2). This, along with strong FDR control, highlights our method’s effectiveness in SRS signal detection.
Figure 4: Heatmap of 20 selected PTs. Cells not detected as signals by our method are colored gray. White-to-blue shades represent the posterior probability of $1-z_{ij}^{l}$, with darker (lighter) blue indicating a higher (lower) probability that the signal strength originates from the global signal strength distribution. The letters “H”, “B”, “G”, and “L” indicate signal detection by the DP Hu et al., BCPNN, GPS, and likelihood ratio test, respectively.
Finally, we focus on inferring structural zeros through the estimated drug-specific zero-inflation probabilities $\omega_{j}$ as returned by our DP ZIP local-global model. For comparison, we also monitor the corresponding estimates provided by the frequentist ZIP (pseudo) LRT method. The two methods yield largely similar results: our approach produces point estimates (posterior mean) for $\omega_{j}$ as 0.02, 0.10, 0.11, 0.08, 0.02, and 0.03 for Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, respectively. The corresponding frequentist maximum likelihood estimates are 0.00, 0.07, 0.16, 0.06, 0.00, and 0.00, respectively.
6 DISCUSSION
We propose nonparametric Bayesian models and their MCMC implementation for SRS data mining, balancing flexibility with shrinkage while capturing between-drug associations. Signal detection is performed through hypothesis tests that control FDR and minimize FNR, both estimated from MCMC posteriors. Our approach achieves significantly higher sensitivity than existing methods, particularly for rarer AEs and/or new drugs.
There are several limitations with our current approach that motivate interesting future directions. First, we focus here primarily on signal detection, i.e., to identify critical AEs of concern for each drug. The clusters of AEs produced by our DP priors have not been investigated but may reveal additional, deeper insights by identifying groups of “similar AEs”. A reformulation via mixed membership models could further enhance interpretability by allowing the same AE to appear in multiple clusters/groups. Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI; Poleksic and Xie, 2019), possibly via appropriate prior distributions, may further enhance signal detection. Finally, our current framework aggregates drug-AE data across time points for a cross-sectional analysis. Extending the model to accommodate temporal variations in SRS data could enable formal identification of evolving AE patterns.
Acknowledgment. The authors are deeply grateful to Prof. Marianthi Markatou at the University at Buffalo for her insightful comments and helpful suggestions on the manuscript.
Supplementary Material
This supplement provides a brief review of existing SRS data mining methods, additional details about the DP local-global models, detailed steps for the MCMC algorithms implementing our models, additional computation details, and additional simulation results. Sections, tables, and figures in this supplement are labeled with a prefix “S.”, such as, Section S.1, Table S.1, and Figure S.1, etc.
Appendix S.1 A REVIEW OF EXISTING BAYESIAN/EMPIRICAL BAYESIAN SRS DATA MINING METHODS AND THE ZIP MODEL-BASED LIKELIHOOD RATIO TEST METHOD
S.1.1 The GPS model
The GPS method (DuMouchel,, 1999) employs an empirical Bayes approach for stabilized signal detection within a Poisson model-based framework for the report counts $n_{ij}$. Specifically, we assume $n_{ij}\sim$ Poisson$(E_{ij}\lambda_{ij})$, with $\lambda_{ij}$ parametrizing the signal strength for the drug-$j$ and AE-$i$ pair. A two-component gamma mixture prior for $\lambda_{ij}$ is considered:
$\lambda_{ij}\sim\kappa\text{Gamma}(\alpha_{1},\beta_{1})+(1-\kappa)\text{Gamma}(\alpha_{2},\beta_{2}),$
where $\kappa\in[0,1]$ is the mixing weight hyperparameter, and $\alpha_{1},\beta_{1}>0$, and $\alpha_{2},\beta_{2}>0$ are mixture component-specific gamma hyperparemeters. Here Gamma($\alpha$, $\beta$) denotes a probability distribution with density $f_{\text{gamma}}(y\mid\alpha,\beta)=\frac{1}{\Gamma(\alpha)}\beta^{\alpha}y^{\alpha-1}e^{-\beta y}$, $y>0$; $\alpha,\beta>0$. The posterior distribution of $\lambda_{ij}$ given $n_{ij}$, $E_{ij}$, and the hyperparameters is:
$\lambda_{ij}\mid n_{ij},E_{ij}\sim\kappa_{ij}\text{Gamma}(\hat{\alpha_{1}}+n_{ij},\hat{\beta_{1}}+E_{ij})+(1-\kappa_{ij})\text{Gamma}(\hat{\alpha_{2}}+n_{ij},\hat{\beta_{2}}+E_{ij})$
where $\kappa_{ij}=\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})/[\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})+(1-\kappa)f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{2},\beta_{2})]$. Under an empirical Bayes approach, the point estimates $\hat{\kappa}$, $\hat{\alpha}_{1}$, $\hat{\beta}_{1}$, $\hat{\alpha}_{2}$, and $\hat{\beta}_{2}$ for the hyperparameters are obtained by maximizing their marginal likelihoods given $n_{ij}$ and $E_{ij}$. The resulting posterior of $\lambda_{ij}$ conditional on these estimated hyperparameters is used for inference and signal detection. Specifically, DuMouchel and Pregibon, (2001) define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds 2.
S.1.2 The BCPNN method
The BCPNN approach (Bate et al.,, 1998) assumes that conditional on the grand total count $n_{\bullet\bullet}$, the cell-specific counts $n_{ij}$, the marginal row totals $n_{i\bullet}$, and the marginal column totals $n_{\bullet j}$ each distributed as binomial-beta: (a) $n_{ij}\mid p_{ij}\sim\text{Binomial}(n_{\bullet\bullet},p_{ij}),\ p_{ij}\sim\text{Beta}(\alpha_{ij},\beta_{ij})$, (b) $n_{i\bullet}\mid p_{i\bullet}\sim\text{Binomial}(n_{\bullet\bullet},p_{i\bullet}),\allowbreak\ p_{i\bullet}\sim\text{Beta}(\alpha_{i\bullet},\beta_{i\bullet})$, (c) $n_{\bullet j}\mid p_{\bullet j}\sim\text{Binomial}(n_{\bullet\bullet},p_{\bullet j}),\allowbreak\ p_{\bullet j}\sim\text{Beta}(\alpha_{\bullet j},\beta_{\bullet j})$, where $p_{ij}$ is the cell reporting rate, and $p_{i\bullet}$ and $p_{\bullet j}$ are the marginal row and column reporting rates. (Note that the binomial model assumptions for $n_{ij}$, $n_{i\bullet}$ and $n_{\bullet j}$ can be understood as consequences of an independent Poisson model for $n_{ij}$ with certain choices of $p_{ij}$, $p_{i\bullet}$, and $p_{\bullet j}$.). Bate et al., (1998) consider the information content $\text{IC}_{ij}=\log_{2}\left[p_{ij}/(p_{i\bullet}p_{\bullet j})\right]$ as a measure of signal strength for the drug-$j$ AE-$i$ pair, with a higher IC indicating a stronger association. The authors suggest an empirical Bayes estimation of these quantities using $\hat{\alpha}_{ij}=\hat{\alpha}_{i\bullet}=\hat{\beta}_{i\bullet}=\hat{\alpha}_{\bullet j}=\hat{\beta}_{\bullet j}=1$ and $\hat{\beta}_{ij}=1/(E(p_{i\bullet}\mid n_{i\bullet})E(p_{\bullet j}\mid n_{\bullet j}))-1$, with the denominator of $\hat{\beta}_{ij}$ being computed using the posterior distributions $p_{i\bullet}\mid n_{i\bullet}\sim\text{Beta}(\hat{\alpha}_{i\bullet}+n_{i\bullet},\hat{\beta}_{i\bullet}+n_{\bullet\bullet}-n_{i\bullet})$ and $p_{\bullet j}\mid n_{\bullet j}\sim\text{Beta}(\hat{\alpha}_{\bullet j}+n_{\bullet j},\hat{\beta}_{\bullet j}+n_{\bullet\bullet}-n_{\bullet j})$. The consequent (empirical Bayes) posterior distribution of $\text{IC}_{ij}$ can be approximated using Monte Carlo simulation or an asymptotic normal distribution. A signal is detected at the drug-$j$ and AE-$i$ pair if the 2.5th posterior percentile for $\text{IC}_{ij}$ exceeds 0.
S.1.3 DP Poisson method of Hu et al., (2015)
Hu et al., (2015) consider the same likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but extend the two-gamma mixture prior for $\lambda_{ij}$ to a DP (infinite) mixture prior:
$\lambda_{ij}\mid(\alpha_{j},M_{0j})\sim\text{DP}(\alpha,M_{0j})=\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}$
where the base distribution $M_{0j}\equiv$ Gamma$(a_{j},a_{j})$ generates $\theta_{hj}\sim M_{0j}$ and the concentration hyperparameter $\alpha>0$ generates the mixture weights $\{\eta_{hj}\}_{h=1}^{\infty}$ satisfying $\sum_{h=1}^{\infty}\eta_{hj}=1$ with
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj}),\ h=2,\dots;\ v_{hj}\sim\text{independent Beta}(1,\alpha)\text{ for }h=1,\dots$ (5)
leveraging a stick-breaking representation (Sethuraman,, 1994) for each $j$. Hu et al., (2015) consider full Bayesian inference on $\lambda_{ij}$ by first eliciting a prior distribution on the hyperparameters $a_{j}$ and $\alpha_{j}$—specifically, independent vague priors $a_{j}\sim\text{Uniform}(0,1)$ and $\alpha_{j}\sim\text{Uniform}(0.2,10)$; $j=1,\dots J$—and then utilizing MCMC sampling from the resulting joint posterior. Using MCMC posterior draws for $\lambda_{ij}$, a signal at the drug-$j$ and AE-$i$ pair is identified if the 5th posterior percentile of $\lambda_{ij}$ exceeds a threshold of 2.
S.1.4 ZIP model based Pseudo LRT (frequentist) method
The likelihood ratio test based method and its variants (Huang et al.,, 2011, 2013, 2014) have an advantage in FDR control while maintaining high sensitivity. Additionally, the use of a post-hoc threshold makes it more suitable for practical applications. Consider the reporting count $n_{ij}\sim\text{ZIP}(n=n_{ij}\mid\lambda_{ij}\times E_{ij},\omega_{j})$, where $\lambda_{ij}$ is the relative reporting rate and $E_{ij}=n_{i\cdot}n_{\cdot j}/n_{\cdot\cdot}$ represents the expected count under independence. The pseudo LRT, as proposed by Chakraborty et al., (2022), is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands. Similar to the standard LRT method, this approach first tests a global alternative hypothesis $H_{aj}:\lambda_{ij}>1$ for at least one $i\in\{1,\dots,I\}$ against the null hypothesis $H_{0j}:\lambda_{ij}=1$ for all $i\in\{1,\dots,I\}$ for a given drug-$j$. If $H_{aj}$ is rejected, a step-down procedure is used to conduct individual tests on $H_{aij}:\lambda_{ij}>1$.
The ordinary LRT methods estimate the nuisance parameter $\omega_{j}$ from $H_{0j}$ and $H_{0j}\cup H_{aj}$ individually, the pseudo LRT is constructed using a global estimate $\hat{\omega}_{j}$ of $\omega_{j}$ consistent under $H_{0j}\cup H_{aj}$ in both the numerator and the denominator of the likelihood ratio test statistic. The resulting statistic is given by
$\text{LR}_{ij}=\exp(-(\hat{\lambda}_{ij}-1)E_{ij})\hat{\lambda}_{ij}^{n}.$
The maximum likelihood ratio statistic for drug-$j$ is then defined as $\text{MLR}_{j}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\}\end{subarray}}\text{LR}_{ij}$. In the global test, a parametric bootstrap can be used to approximate the p-value $P_{j}=\Pr(\text{MLR}_{j}\geq\text{MLR}_{j}^{\text{obs}})$ using (parametric bootstrap) draws for $\text{LR}_{ij}$ under $H_{0j}$ and hence $\text{MLR}_{j}$, where $\text{MLR}_{j}^{\text{obs}}$ is the observed test statistic. If the global test is rejected, the step-down p-values can be computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$. For multiple drugs, the same approach can be applied, defining an overall maximum likelihood ratio as $\text{MLR}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\},j\in\{1,\dots,J\}\end{subarray}}\text{LR}_{ij}$. This step-down procedure avoids $\alpha$-inflation by using the same global test statistic, thus controlling the false discovery rate at the $\alpha$ level.
A consistent estimate of $\omega_{j}$ is obtained by maximizing the profile likelihood $\ell_{j}(\omega_{j})$, with $\hat{\lambda}_{ij}=\max\{1,n/E_{ij}\}$ as the MLE of $\lambda_{ij}$. The inference on zero-inflation is based on the estimate $\hat{\eta}_{ij}(\lambda_{ij},\hat{\omega}_{j})=\frac{\hat{\omega}_{j}}{\hat{\omega}_{j}+(1-\hat{\omega}_{j})\exp(-\lambda_{ij}E_{ij})}$, representing the estimated probability of zero-inflation for cells with zero count. A prediction of the zero-inflation indicator $z_{ij}$ can be made by generating Bernoulli samples from $\hat{\eta}_{ij}$.
Appendix S.2 ADDITIONAL DETAILS FOR THE DP LOCAL-GLOBAL FRAMEWORK
In this section, we generalize the model to encompass different variants depending on the direction of information sharing. Moreover, we provide additional details of the DP local-global framework. And we propose a simpler shrinkage model as an alternative for the “Other drugs” as described in Section 3.2.
S.2.1 Complete hierarchical Bayesian representation of our models
In Section 3.2 in the main text, we present the concept of the DP Poisson local-global model. The fully formulated hierarchical representations are provided in the following two sections for the DP Poisson and the DP ZIP local-global model.
S.2.1.1 DP Poison local-global model
A hierarchical representation of the complete model, including vague priors for the hyperparameters, is presented as follows.
likelihood:
$n_{ij}\sim\text{Poisson}(\lambda_{ij}E_{ij})$
[…]
$\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
$\lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D$
$z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
DP local prior for drug-$j$:
$D_{j}\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}\approx\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}\sim\text{DP}(\alpha_{j},G(\beta_{j}))$
$\theta_{hj}\sim G(\beta_{j})\equiv\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj}),h\geq 2$
$v_{hj}\sim\text{Beta}(1,\alpha_{j}),h=1,\dots,K-1;\ v_{Kj}=1$
DP global prior:
$D\equiv\sum_{h=1}^{\infty}\eta_{h}\delta_{\{\theta_{h}\}}\approx\sum_{h=1}^{K}\eta_{h}\delta_{\{\theta_{h}\}}\sim DP(\alpha,G(\beta))$
$\theta_{h}\sim G(\beta)\equiv\text{Gamma}(\text{shape}=\beta,\text{rate}=\beta)$
$\eta_{1}=v_{1},\ \eta_{h}=v_{h}\prod_{t=1}^{h-1}(1-v_{t}),h\geq 2$
$v_{h}\sim\text{Beta}(1,\alpha)$
Gamma shrinkage prior for a baseline drug-$J$:
$\lambda_{iJ}\sim\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$
vague proper priors for hyperparameters:
$\pi^{g}\sim\text{Uniform}(0,1)$
$\alpha_{j},\alpha\sim\text{Gamma}(\text{shape}=1,\text{rate}=3)$
$\beta_{j},\beta\sim\text{Cauchy}^{+}(0,0.5)$
$\tau\sim\text{Cauchy}^{+}(0,0.5)$
S.2.1.2 DP ZIP local-global model
A hierarchical representation of the complete model, including vague priors for the hyperparameters, is presented as follows.
likelihood:
$n_{ij}\sim y\text{Poisson}(\lambda_{ij}E_{ij})+(1-y)\delta_{\{0\}}$
$y\sim\text{Bernoulli}(\omega_{j})$
[…]
$\lambda_{ij}\sim z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$
$\lambda_{ij}^{l}\sim D_{j};\ \lambda_{i}^{g}\sim D$
$z_{ij}^{l}\sim\text{Bernoulli}(\pi^{l})$
DP local prior for drug-$j$:
$D_{j}\equiv DP(\alpha_{j},G(\beta_{j}))\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}\approx\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$
$\theta_{hj}\sim G_{0j}\equiv\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj}),h\geq 2$
$v_{hj}\sim\text{Beta}(1,\alpha_{j});v_{Kj}=1$
DP global prior:
$D\equiv DP(\alpha,G(\beta))\equiv\sum_{h=1}^{\infty}\eta_{h}\delta_{\{\theta_{h}\}}\approx\sum_{h=1}^{K}\eta_{h}\delta_{\{\theta_{h}\}}$
$\theta_{h}\sim G_{0}\equiv\text{Gamma}(\beta,\beta)$
$\eta_{1}=v_{1},\ \eta_{h}=v_{h}\prod_{t=1}^{h-1}(1-v_{t}),h\geq 2$
$v_{h}\sim\text{Beta}(1,\alpha)$
Gamma shrinkage prior for a baseline drug-$J$:
$\lambda_{iJ}\sim\text{Gamma}(\text{shape}=\tau,\text{rate}=\tau)$
vague proper priors for hyperparameters:
$\pi^{g}\sim\text{Uniform}(0,1)$
$\alpha_{j},\alpha\sim\text{Gamma}(\text{shape}=1,\text{rate}=3)$
$\beta_{j},\beta\sim\text{Cauchy}^{+}(0,0.5)$
$\tau\sim\text{Cauchy}^{+}(0,0.5)$
S.2.2 Some notes on the “global-only” model
A special case of our local-global models leads to a reduced “global-only” model when $\pi^{g}=0$, i.e., $\lambda_{ij}=\lambda_{i}^{g}$ for $j=1,\dots,J-1$. Under this model there is no provision for different $\lambda_{ij}$ for any $j\neq J$. The global model has a simplified likelihood $n_{ij}\sim\text{Poisson}(\lambda_{i}E_{ij})$ here $\lambda_{i}^{g}\sim D$ if $j\neq J$, which produces $\sum_{j=1}^{J-1}n_{ij}\sim\text{Poisson}(\lambda_{i}\sum_{j=1}^{J-1}E_{ij})$. The model reduction also impacts the signal detection. For the local-global framework, hypothesis testing is conducted for each individual $\lambda_{ij}$. On the other hand, under the global model, the test is simplified to evaluating whether $\lambda_{i}^{g}>1$. Testing $\lambda_{i}^{g}>1$ entails assessing if the summarized Poisson count $\sum_{i}n_{ij}$ is large compared to the expected count $\sum_{i}E_{ij}$, owing to the reduced model. This approach represents a “global test”, whereas our proposed local-global model allows conducting hypothesis testing for each individual drug-AE pair. As a result, directly comparing the global model with the local-global model is not feasible, as they target different types of assessment.
Appendix S.3 COMPUTATIONAL DETAILS
In this section, we first provide MCMC algorithms implementing our models in Section S.3.1. The MCMC algorithms are derived using finite mixture approximation of the underlying DP processes (Ishwaran and Zarepour,, 2002; Frühwirth-Schnatter and Malsiner-Walli,, 2019), and combine Gibbs sampling updates with stepping-out slice sampling updates (Neal,, 2003) for various parameters. Some additional notes on the MCMC implementation, including initial value generation for our MCMC algorithms are then provided in Section S.3.2. Next we provide detailed notes on the grid-search method identifying optimal quantile-level $p$ and threshold $k$ used for our hypothesis tests for signal detection (Section S.3.3). The JAGS code (Plummer et al.,, 2003) for Hu et al., (2015)’s models is shown as a reference in Section S.1.3. The example statin dataset is included in Section S.3.5.
In our notation below, $\text{DPSB}_{K}(\bm{\eta}=(\eta_{1},\dots,\eta_{K}),\bm{\theta}=(\theta_{1},\dots,\theta_{K}),\alpha,\beta)$ denotes parameters associated with the model $\sum_{h=1}^{K}\eta_{h}1_{\{\theta_{h}\}}\sim\text{DP}(\alpha,G(\beta))$, comprising the stick-breaking weights $\bm{\eta}=(\eta_{1},\dots,\eta_{K})$ and atoms $\bm{\theta}=(\theta_{1},\dots,\theta_{K})$ and DP concentration hyperparamter $\alpha$ and base (gamma) distribution hyperparameter $\beta$.
S.3.1 MCMC algorithms
Algorithm S.1 One Iteration of the MCMC algorithm updating $\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$ parameters from its posterior given data $(n_{i},E_{i},z_{i})$ under the model: $n_{i}\sim\text{Pois}(\lambda_{i}E_{i})$, $\lambda_{i}\sim\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$, $\alpha\sim\text{Exponential}(\text{rate}=\psi_{\alpha})$, $\beta\sim\pi_{\beta}(\beta)$, and $z_{i}$ indicating if observation $i$ contributes to the model ($z_{i}=1$) or not ($z_{i}=0$); $i=1,\dots,N$.
  1. 1.
    Generate allocations $\bm{S}=(S_{1},\dots,S_{N})$ using
    $P(S_{i}=h\mid\bm{\eta},\bm{\theta},\bm{z},\bm{n},\bm{E})=\frac{\eta_{h}\{f_{\text{Poisson}}(n_{i}\mid\theta_{h}E_{i})\}^{z_{i}}}{\sum_{h^{\prime}=1}^{K}\eta_{h^{\prime}}\{f_{\text{Poisson}}(n_{i}\mid\theta_{h^{\prime}}E_{i})\}^{z_{i}}}$
    for $h=1,\dots,K$. Here $f_{\text{Poisson}}(n\mid\theta)$ denotes the probability mass function of the Poisson($\theta$) distribution evaluated at $n$.
  2. 2.
    Generate $\theta_{1},\dots,\theta_{K}$ independently from
    $\theta_{h}\mid\bm{S},\alpha,\bm{n},\bm{E},\bm{z}\sim\text{Gamma}\!\Biggl{(}\text{shape}=\sum_{i=1}^{N}n_{i}z_{i}1(S_{i}=h)+\alpha,\text{rate}=\sum_{i=1}^{N}E_{i}z_{i}1(S_{i}=h)+\alpha\Biggr{)}$
    1. 2.1.
      Calculate $\lambda_{i}=\theta_{S_{i}}$, $i=1,\dots,N$
  3. 3.
    Generate $\beta$ from the density
    $\pi(\beta\mid\bm{\theta})\propto\prod_{h=1}^{K}f_{\text{Gamma}}(\theta_{h}\mid\text{shape}=\beta,\text{rate}=\beta)\pi_{\beta}(\beta)$
    Here $f_{\text{Gamma}}(\theta\mid\text{shape}=\beta,\text{rate}=\beta)$ denotes the probability density function of the Gamma($\beta,\beta$) distribution at $\theta$. $\pi(\beta)$ is the prior distribution of $\beta$. Direct random sampling from this density is not feasible. Instead, we use stepping-out slice sampling (Neal,, 2003) to generate Markov chain draws for $\beta$.
  4. 4.
    Generate $v_{1},\dots,v_{K-1}$ independently from:
    $v_{h}\mid\bm{S},\alpha\sim\text{Beta}\left(1+\sum_{i=1}^{N}1(S_{i}=k),\alpha+\sum_{i=1}^{N}1(S_{i}>k)\right).$
    Then set $v_{K}=1$ and calculate $\eta_{1},\dots,\eta_{K}$ from $v_{1},\dots,v_{K}$ using
    $\eta_{1}=v_{1},\quad\eta_{h}=v_{h}\prod_{t=1}^{h-1}(1-v_{t}),\quad h=2,\dots,K.$
  5. 5.
    Generate $\alpha$ from
    $\alpha\mid v_{1},\dots,v_{K}\sim\text{Gamma}\left(\text{shape}=K,\text{rate}=-\sum_{h=1}^{K-1}\log(1-v_{h})+\psi_{\alpha}\right).$
Algorithm S.2 One Step Iteration of DP Poisson local-global column model (Given data $\{(n_{ij},E_{ij},z_{ij}^{l}):i=1,\dots,I;j=1,\dots,J\}$ )
  1. 1.
    Generate global $\text{DPSB}_{K}(\bm{\eta},\bm{\theta},\alpha,\beta)$ parameters using Algorithm S.1, with $\{(\tilde{n}_{i},\tilde{E}_{i},1(\sum_{j=1}^{J-1}z_{ij}^{l}<J-1)):i=1,\dots,I\}$ as data where
    $\tilde{n}_{j}=\sum_{j=1}^{J-1}n_{ij}(1-z_{ij}^{l}),\quad\tilde{E}_{j}=\sum_{j=1}^{J-1}E_{ij}(1-z_{ij}^{l}).$
    1. 1.1.
      Set $\lambda_{i}^{\text{g}}=\lambda_{i}$ returned from Algorithm S.1.
  2. 2.
    For $j=1,\dots,J-1$, generate local $\text{DPSB}(\bm{\eta}_{j},\bm{\theta}_{j},\alpha_{j},\beta_{j})$ using Algorithm S.1 using $\{(n_{ij},E_{ij},z_{ij});i=1,\dots,I\}$ as data.
    1. 2.1.j
      For this $j$, set $\lambda_{ij}^{l}=\lambda_{i}$ returned from Algorithm S.1.
  3. 3.
    Generate “local” allocations:
    $z_{ij}^{l}\mid n_{ij},E_{ij},z_{ij}^{l},\lambda_{ij}^{l},\lambda_{i}^{g}\sim\text{Bernoulli}(p_{ij}^{l}),$
    where
    $p_{ij}^{l}=\frac{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi^{l})f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
  4. 4.
    Define $\lambda_{ij}=z_{ij}^{l}\lambda_{ij}^{l}+(1-z_{ij}^{l})\lambda_{i}^{g}$ for all $i=1,\dots,I$ and $j=1,\dots,(J-1)$.
  5. 5.
    Generate $\pi^{l}$ using:
    $\pi^{l}\mid z_{ij}^{l}\sim\text{Beta}\left(a^{l}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}z_{ij},b^{l}+\sum_{i=1}^{I}\sum_{j=1}^{J-1}(1-z_{ij})\right),$
    Notes: $a^{l}$ = 1, $b^{l}$ = 1
  6. 6.
    Generate $\lambda_{1J},\dots,\lambda_{IJ}$ from:
    $\lambda_{iJ}\mid\tau,n_{iJ},E_{iJ}\sim\text{Gamma}(\text{shape}=n_{iJ}+\tau,\text{rate}=E_{iJ}+\tau).$
  7. 7.
    Generate $\beta_{J}$ from:
    $\pi(\beta_{J}\mid\lambda_{1J},\dots,\lambda_{IJ})\propto\prod_{i=1}^{I}f_{\text{Gamma}}(\lambda_{ij}\mid\text{shape}=\tau,\text{rate}=\tau)\pi_{\tau}(\tau).$
Algorithm S.3 One Iteration of DP ZIP local-global Model Given data $(n_{ij},E_{ij},z_{ij}^{l})$
  1. 1.
    Generate the ZI indicator $y_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, where:
    $y_{ij}\mid\omega_{j},\lambda_{ij},n_{ij},E_{ij}\sim\text{Bernoulli}(p_{ij}^{\text{ZI}}),$
    where
    $p_{ij}^{\text{ZI}}=\begin{cases}\frac{\omega_{j}}{\omega_{j}+(1-\omega_{j})\exp(-\lambda_{ij}E_{ij})},&\text{if }n_{ij}=0,\\ 0,&\text{if }n_{ij}\geq 1.\end{cases}$
  2. 2.
    Generate $\omega_{1},\dots,\omega_{J}$ independently with:
    $\omega_{j}\mid y_{1j}\dots y_{Ij}\sim\text{Beta}\left(1+\sum_{i=1}^{I}y_{ij},1+\sum_{i=1}^{I}(1-y_{ij})\right).$
  3. 3.
    Run Algorithm S.2 with all $(n_{ij},E_{ij})$ such that $y_{ij}=0$.
S.3.2 Initial value generation for the MCMC algorithms
Given the observed data $n_{ij}$ and $E_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, and a prespecified number $K$ of DP mixture components for finite approximation, we initialize the number of nonempty components $K_{+}$ to the smallest integer greater than $\min\{K/2,I-1\}$. The local-global indicators $z_{ij}^{l}$ are randomly initialized by independently sampling each cell $(i,j)$ from a $\text{Bernoulli}(0.5)$ distribution. The DP parameters for each local and global DP process are then initialized using the subset of data $(n_{ij},E_{ij})$ assigned to that process, as determined by the indicators $z_{ij}^{l}$, for $i=1,\dots,J-1$.
For the initialization of parameters in the local DP prior for the $j$-th drug ($j=1,\dots,J-1$), we perform $k$-means clustering with the initialized $K_{+}$ number of clusters on the set ${(n_{ij}+0.5)/(E_{ij}+0.5):z_{ij}=1,\ i=1,\dots,I}$, and use the resulting cluster centers as initial values for the atoms $\theta_{hj}$, $h=1,\dots,K_{+}$, of the non-empty clusters. The probability weights $\eta_{hj}$ for these non-empty clusters are initialized using the relative frequencies of the $k$-means clusters, and the clusters are then relabeled to ensure $\eta_{1j}\geq\eta_{2j}\geq\dots\geq\eta_{K_{+}j}$. These atoms $\theta_{hj}$ are then used for numerical maximum likelihood estimation of $\beta$ with the model $\theta_{1j},\dots,\theta_{K_{+},j}\sim\text{independent~{}Gamma}(\text{shape}=\beta,\text{rate}=\beta)$, and the resulting estimate $\hat{\beta}$ is used to initialize $\beta_{j}$. The atoms $\theta_{h^{\prime}j}$ for the remaining empty clusters ($h^{\prime}=K_{+}+1,\dots,K$) are randomly drawn from $\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$ using the initialized $\beta_{j}$. Their corresponding weights $\eta_{h^{\prime}j}$ are initialized to $(\min_{h^{\prime}\in{1,\dots,K_{+}}}\eta_{h^{\prime}j})/10$, and all weights $\eta_{hj}$ for $h=1,\dots,K$ are subsequently renormalized to ensure $\sum_{h=1}^{K}\eta_{hj}=1$. The local signal strengths $\lambda_{ij}^{l}$ are then initialized for each $i$ using the cluster center (atom) corresponding to the $k$-means cluster containing point $i$ if $z_{ij}=1$, or by random sampling from ${\theta_{1j},\dots,\theta_{Kj}}$ with probabilities ${\eta_{1j},\dots,\eta_{Kj}}$. Finally, the concentration hyperparameter $\alpha_{j}$ is initialized by random sampling from a $\text{Gamma}(\text{shape}=1,\text{rate}=2)$ distribution.
The parameters for the global DP prior are analogously initialized, but using the aggregated observed and expected counts $\{\tilde{n}_{i}=\sum_{j=1}^{J-1}n_{ij}(1-z_{ij}),\tilde{E}_{i}=\sum_{j=1}^{J-1}E_{ij}(1-z_{ij})\}$ as the data with all AEs $i$ such that $\sum_{j=1}^{J-1}(1-z_{ij})>0$.
Finally, the reference drug $J$ shrinkage hyperparameter $\tau$ is initialized at $1$, while the zero-inflation parameters $\omega_{j}$ (if included in the model) are initialized each at $0.01$.
S.3.3 Optimization method in pFDR calculation
A sequence of threshold values for $k$, ranging from 1.1 to 3 in small increments, is established. For our computations we specified this grid to be {1.1, 1.11, 1.12, 1.13, 1.14, 1.16, 1.17, 1.18, 1.19, 1.2, 1.25, 1.28, 1.31, 1.33, 1.36, 1.39, 1.42, 1.44, 1.47, 1.5, 1.6, 1.76, 1.91, 2.07, 2.22, 2.38, 2.53, 2.69, 2.84, 3}. Similarly, a set of distinct small quantile levels for $p$ is chosen ({0.01, 0.015, 0.02, 0.025, 0.03, 0.035, 0.04, 0.045, 0.05, 0.055, 0.06, 0.065, 0.07, 0.075, 0.08, 0.085, 0.09, 0.095, 0.1} in our computations). These define a range of possible combinations $(p,k)$ to be considered.
For every quantile level in the $p$-grid, we obtain a matrix of test statistics $T_{ij}(p)$ based on the $p$-th posterior quantiles for $\lambda_{ij}$ computed from their MCMC draws. Subsequently, we compute FDR$(p,k)$ and FNR$(p,k)$ estimates using the formulas presented in the main text (Section 3.4) for all possible $p$ and $k$ defined by the grid. From the set of all $(p,k)$ pairs, we then retain only those pairs with computed FDR $\leq\alpha$ for a preset threshold $\alpha$ ($=0.05$ in our analyses). Finally, among the retained $(p,k)$, we identify the pair that minimizes FNR. If multiple $(p,k)$ yield the same minimum FNR, we select the pair with the smallest $p$ and the largest $k$.
S.3.4 JAGS code for DP Hu et al. model
1 # Likelihood
2 for (i in 1:I) {
3 for (j in 1:J) {
4 n[i, j] ~ dpois(E[i, j] * lambda[i, j])
5 }
6 }
7
8 # Prior
9 for (j in 1:J) {
10 alpha[j] ~ dunif(0.2, 10)
11 a[j] ~ dunif(0, 1)
12
13 v[j, 1] ~ dbeta(1, alpha[j]); T(0.00001,0.99999) # to ensure in (0, 1)
14 w[j, 1] <- v[j, 1]
15
16 for (h in 2:K) {
17 v[j, h] ~ dbeta(1, alpha[j]); T(0.00001,0.99999) # to ensure in (0, 1)
18 w[j, h] <- v[j, h] * prod( (1 - c(v[j, 1:(h-1)])) )
19 }
20
21 for (h in 1:K) {
22 theta[j, h] ~ dgamma(a[j], a[j]); T(0.000001, ) # to ensure > 0
23 }
24
25 for (i in 1:I) {
26 ind[i, j] ~ dcat(c(w[j,]))
27 lambda[i, j] <- theta[j, ind[i, j]]
28 }
29 }
Listing 1: JAGS exmpale
S.3.5 Example statin dataset
We present the reference SRS contingency table for statin drug in Table S.1. It is used for data generation in Section 4 and data analysis in Section 5.
Table S.1: The SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases.
Adverse Event Atorvastatin Fluvastatin Lovastatin Pravastatin Rosuvastatin Simvastatin Other
Acute Kidney Injury 1353 42 7 154 689 823 355651
Anuria 71 0 0 2 43 62 10403
Blood Calcium Decreased 14 2 0 0 110 17 15918
Blood Creatine Phosphokinase Abnormal 34 0 0 0 8 11 261
Blood Creatine Phosphokinase Increased 1175 125 32 200 562 768 23805
Blood Creatine Phosphokinase Mm Increased 2 0 0 0 9 0 14
Blood Creatinine Abnormal 27 0 0 0 5 3 3385
Blood Creatinine Increased 227 10 0 17 210 97 74742
Chromaturia 340 10 6 33 174 114 19294
Chronic Kidney Disease 152 16 2 19 177 37 339179
Compartment Syndrome 53 0 0 1 12 12 2644
Creatinine Renal Clearance Decreased 6 0 0 2 124 6 7768
Diaphragm Muscle Weakness 14 0 0 0 8 1 94
Electromyogram Abnormal 2 0 0 2 0 0 132
End Stage Renal Disease 30 0 0 0 19 6 97553
Glomerular Filtration Rate Abnormal 8 0 0 0 1 0 1069
Glomerular Filtration Rate Decreased 59 1 0 8 39 29 13190
Hypercreatininaemia 0 0 0 0 8 0 648
Hypocalcaemia 36 0 0 16 8 18 23102
Muscle Disorder 291 2 7 21 191 87 7329
Muscle Enzyme Increased 48 1 0 0 13 9 410
Muscle Fatigue 85 0 2 16 30 39 4257
Muscle Haemorrhage 24 0 0 5 13 4 3806
Muscle Necrosis 68 2 0 1 10 20 662
Muscle Rupture 181 25 0 61 36 120 3219
Muscular Weakness 1857 45 31 152 808 859 111003
Musculoskeletal Discomfort 137 18 15 25 187 93 19931
Musculoskeletal Disorder 56 3 0 9 65 73 25881
Musculoskeletal Pain 420 3 2 38 324 228 82576
Myalgia 5362 341 151 939 2757 3216 143819
Myasthenic Syndrome 1 0 9 0 3 6 643
Myoglobin Blood Increased 71 4 0 4 16 39 1003
Myoglobin Blood Present 2 0 0 0 0 0 8
Myoglobin Urine Present 0 0 0 0 1 2 70
Myoglobinaemia 15 0 0 0 0 0 62
Myoglobinuria 26 4 0 1 1 9 296
Myopathy 849 64 45 145 377 544 6695
Myopathy Toxic 31 0 0 1 2 21 457
Myositis 219 8 10 28 62 141 7482
Necrotising Myositis 279 0 0 2 10 52 278
Oliguria 52 0 0 4 24 37 7590
Renal Failure 534 26 11 69 225 195 250710
Renal Impairment 390 52 11 37 161 169 103343
Renal Tubular Necrosis 40 0 0 12 10 24 12762
Rhabdomyolysis 2041 52 44 163 936 1376 31707
Tendon Discomfort 9 0 0 3 22 10 794
Other Pt 180699 4886 2845 20296 113960 76068 61724222
Appendix S.4 ADDITIONAL SIMULATION RESULTS
We present detailed values of the simulation to compare among our drug (column)-based models (Poisson/ZIP), DP Hu et al., and posterior quantile optimized DP Hu et al. method. Four evaluation metrics (FDR, sensitivity, average type I error, and F-score) are evaluated as well as the main text with same simulation settings in Table 1. Frequentist and Bayesian MSE are compared to further evaluate the loss of the estimation.
S.4.1 Simulation results for DP-based models
We present the four evaluation metrics in Tables S.2, S.3, S.4, and S.5 to provide a clearer numerical comparison beyond Figure 3. Additionally, we further explored the effect of posterior quantile optimization on the DP Hu et al. in the process of FDR control and minimizing FNR. This allows for a more comprehensive understand the strategy of optimizing the posterior quantile of the test statistic and deterministic threshold simultaneously.
We found that in the DP Hu et al. method, simultaneously optimizing the posterior quantile results in only a slight increase in FDR, which remains well-controlled within a very small level (around 0.01). In return, this optimization leads to an improvement in sensitivity, benefiting all cases. When comparing across all DP-based models, the sensitivity gain improvement brought by incorporating the local-global structure is more significant than from this optimization.
Table S.2: Simulation of FDR for DP-based models
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0035 0.0022 0.0050 0.0102
1.2 0.0072 0.0063 0.0056 0.0171
1.3 0.0081 0.0088 0.0053 0.0172
1.4 0.0074 0.0077 0.0061 0.0168
1.5 0.0081 0.0079 0.0059 0.0149
1.7 0.0057 0.0057 0.0050 0.0133
1.9 0.0049 0.0047 0.0039 0.0108
2.0 0.0053 0.0056 0.0046 0.0103
2.5 0.0045 0.0046 0.0043 0.0106
3.0 0.0031 0.0029 0.0034 0.0086
Case 1b Weak correlation ZI 1.1 0.0010 0.0017 0.0060 0.0197
1.2 0.0044 0.0034 0.0096 0.0214
1.3 0.0067 0.0070 0.0070 0.0156
1.4 0.0086 0.0085 0.0051 0.0157
1.5 0.0089 0.0092 0.0059 0.0138
1.7 0.0067 0.0065 0.0056 0.0136
1.9 0.0058 0.0057 0.0048 0.0117
2.0 0.0049 0.0050 0.0045 0.0118
2.5 0.0032 0.0034 0.0032 0.0089
3.0 0.0036 0.0037 0.0042 0.0095
Case 2a Moderate correlation No ZI 1.1 0.0000 0.0010 0.0065 0.0193
1.2 0.0107 0.0128 0.0063 0.0150
1.3 0.0114 0.0132 0.0069 0.0160
1.4 0.0120 0.0134 0.0060 0.0146
1.5 0.0095 0.0098 0.0069 0.0157
1.7 0.0072 0.0079 0.0059 0.0145
1.9 0.0052 0.0056 0.0058 0.0128
2.0 0.0042 0.0045 0.0048 0.0125
2.5 0.0035 0.0035 0.0041 0.0102
3.0 0.0033 0.0032 0.0032 0.0093
Case 2b Moderate correlation ZI 1.1 0.0000 0.0000 0.0055 0.0122
1.2 0.0093 0.0106 0.0065 0.0186
1.3 0.0117 0.0130 0.0069 0.0168
1.4 0.0102 0.0114 0.0061 0.0153
1.5 0.0066 0.0076 0.0054 0.0149
1.7 0.0053 0.0057 0.0058 0.0134
1.9 0.0039 0.0040 0.0059 0.0126
2.0 0.0047 0.0050 0.0045 0.0110
2.5 0.0039 0.0037 0.0042 0.0103
3.0 0.0030 0.0032 0.0041 0.0094
Case 3a Strong correlation No ZI 1.1 0.0000 0.0000 0.0073 0.0133
1.2 0.0084 0.0114 0.0074 0.0183
1.3 0.0073 0.0084 0.0080 0.0184
1.4 0.0086 0.0096 0.0069 0.0145
1.5 0.0061 0.0069 0.0067 0.0150
1.7 0.0045 0.0050 0.0066 0.0136
1.9 0.0045 0.0050 0.0051 0.0129
2.0 0.0039 0.0042 0.0049 0.0118
2.5 0.0033 0.0035 0.0037 0.0096
3.0 0.0030 0.0029 0.0031 0.0085
Case 3b Strong correlation ZI 1.1 0.0000 0.0000 0.0040 0.0171
1.2 0.0106 0.0120 0.0084 0.0162
1.3 0.0077 0.0090 0.0063 0.0161
1.4 0.0078 0.0083 0.0061 0.0153
1.5 0.0068 0.0073 0.0051 0.0125
1.7 0.0044 0.0049 0.0049 0.0119
1.9 0.0030 0.0032 0.0050 0.0112
2.0 0.0035 0.0038 0.0052 0.0113
2.5 0.0033 0.0034 0.0040 0.0092
3.0 0.0036 0.0037 0.0039 0.0094
Table S.3: Simulation of sensitivity for DP-based models
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0016 0.0014 0.0009 0.0021
1.2 0.0427 0.0422 0.0346 0.0435
1.3 0.1227 0.1243 0.0899 0.1041
1.4 0.1923 0.1942 0.1432 0.1606
1.5 0.2444 0.2472 0.1915 0.2115
1.7 0.3153 0.3180 0.2685 0.2897
1.9 0.3692 0.3707 0.3258 0.3475
2.0 0.3926 0.3957 0.3495 0.3720
2.5 0.4706 0.4729 0.4346 0.4554
3.0 0.5251 0.5273 0.4920 0.5128
Case 1b Weak correlation ZI 1.1 0.0012 0.0012 0.0011 0.0022
1.2 0.0406 0.0399 0.0331 0.0416
1.3 0.1216 0.1232 0.0848 0.1006
1.4 0.1948 0.1968 0.1401 0.1581
1.5 0.2426 0.2451 0.1888 0.2102
1.7 0.3182 0.3196 0.2675 0.2898
1.9 0.3709 0.3733 0.3261 0.3457
2.0 0.3921 0.3939 0.3467 0.3678
2.5 0.4717 0.4742 0.4327 0.4564
3.0 0.5227 0.5253 0.4910 0.5124
Case 2a Moderate correlation No ZI 1.1 0.0061 0.0060 0.0012 0.0026
1.2 0.2096 0.2170 0.0526 0.0674
1.3 0.3304 0.3377 0.1334 0.1530
1.4 0.3991 0.4036 0.1998 0.2210
1.5 0.4489 0.4537 0.2499 0.2751
1.7 0.5172 0.5197 0.3359 0.3637
1.9 0.5663 0.5681 0.4043 0.4323
2.0 0.5837 0.5847 0.4310 0.4589
2.5 0.6396 0.6411 0.5363 0.5608
3.0 0.6734 0.6746 0.5984 0.6214
Case 2b Moderate correlation ZI 1.1 0.0067 0.0072 0.0011 0.0025
1.2 0.2067 0.2133 0.0523 0.0659
1.3 0.3309 0.3367 0.1348 0.1535
1.4 0.3968 0.4021 0.1973 0.2186
1.5 0.4438 0.4477 0.2475 0.2723
1.7 0.5138 0.5157 0.3340 0.3599
1.9 0.5601 0.5620 0.4000 0.4274
2.0 0.5772 0.5790 0.4297 0.4570
2.5 0.6329 0.6339 0.5313 0.5561
3.0 0.6696 0.6708 0.5967 0.6205
Case 3a Strong correlation No ZI 1.1 0.0065 0.0076 0.0009 0.0020
1.2 0.2680 0.2784 0.0329 0.0455
1.3 0.4595 0.4682 0.1088 0.1313
1.4 0.5675 0.5746 0.1807 0.2100
1.5 0.6272 0.6330 0.2480 0.2791
1.7 0.7015 0.7049 0.3546 0.3868
1.9 0.7460 0.7482 0.4300 0.4612
2.0 0.7629 0.7644 0.4608 0.4940
2.5 0.8096 0.8106 0.5793 0.6094
3.0 0.8330 0.8336 0.6530 0.6799
Case 3b Strong correlation ZI 1.1 0.0083 0.0094 0.0010 0.0021
1.2 0.2627 0.2769 0.0338 0.0471
1.3 0.4588 0.4688 0.1098 0.1326
1.4 0.5638 0.5717 0.1826 0.2115
1.5 0.6275 0.6333 0.2492 0.2801
1.7 0.7044 0.7080 0.3534 0.3848
1.9 0.7485 0.7507 0.4276 0.4599
2.0 0.7630 0.7651 0.4574 0.4893
2.5 0.8100 0.8109 0.5787 0.6097
3.0 0.8341 0.8348 0.6556 0.6825
Table S.4: Simulation of average type I error for DP-based models
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0001 0.0001 0.0001 0.0002
1.3 0.0002 0.0003 0.0001 0.0003
1.4 0.0002 0.0002 0.0001 0.0004
1.5 0.0003 0.0003 0.0002 0.0005
1.7 0.0002 0.0003 0.0002 0.0005
1.9 0.0002 0.0002 0.0002 0.0005
2.0 0.0003 0.0003 0.0002 0.0005
2.5 0.0003 0.0003 0.0002 0.0006
3.0 0.0002 0.0002 0.0002 0.0006
Case 1b Weak correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0001 0.0001 0.0001 0.0002
1.3 0.0002 0.0002 0.0001 0.0002
1.4 0.0003 0.0003 0.0001 0.0004
1.5 0.0003 0.0004 0.0002 0.0004
1.7 0.0003 0.0003 0.0002 0.0005
1.9 0.0003 0.0003 0.0002 0.0005
2.0 0.0002 0.0002 0.0002 0.0006
2.5 0.0002 0.0002 0.0002 0.0005
3.0 0.0003 0.0003 0.0003 0.0006
Case 2a Moderate correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0005 0.0007 0.0001 0.0002
1.3 0.0006 0.0008 0.0002 0.0004
1.4 0.0008 0.0009 0.0002 0.0005
1.5 0.0006 0.0007 0.0002 0.0006
1.7 0.0005 0.0006 0.0003 0.0007
1.9 0.0004 0.0004 0.0003 0.0007
2.0 0.0003 0.0004 0.0003 0.0008
2.5 0.0003 0.0003 0.0003 0.0007
3.0 0.0003 0.0003 0.0002 0.0007
Case 2b Moderate correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0005 0.0006 0.0001 0.0002
1.3 0.0007 0.0007 0.0001 0.0004
1.4 0.0006 0.0007 0.0002 0.0005
1.5 0.0004 0.0005 0.0002 0.0006
1.7 0.0004 0.0004 0.0003 0.0006
1.9 0.0003 0.0003 0.0003 0.0007
2.0 0.0004 0.0004 0.0003 0.0007
2.5 0.0003 0.0003 0.0003 0.0007
3.0 0.0003 0.0003 0.0003 0.0008
Case 3a Strong correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0005 0.0007 0.0001 0.0002
1.3 0.0006 0.0007 0.0001 0.0004
1.4 0.0009 0.0010 0.0002 0.0005
1.5 0.0007 0.0007 0.0003 0.0007
1.7 0.0005 0.0006 0.0004 0.0008
1.9 0.0005 0.0006 0.0003 0.0009
2.0 0.0005 0.0005 0.0003 0.0009
2.5 0.0004 0.0004 0.0003 0.0009
3.0 0.0004 0.0004 0.0003 0.0009
Case 3b Strong correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0006 0.0007 0.0001 0.0002
1.3 0.0007 0.0008 0.0001 0.0004
1.4 0.0008 0.0009 0.0002 0.0005
1.5 0.0007 0.0008 0.0002 0.0006
1.7 0.0005 0.0006 0.0003 0.0007
1.9 0.0004 0.0004 0.0003 0.0008
2.0 0.0004 0.0005 0.0004 0.0009
2.5 0.0004 0.0004 0.0004 0.0009
3.0 0.0005 0.0005 0.0004 0.0010
Table S.5: Simulation of F-score for DP-based models
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0030 0.0025 0.0018 0.0040
1.2 0.0783 0.0772 0.0650 0.0810
1.3 0.2129 0.2152 0.1615 0.1844
1.4 0.3180 0.3204 0.2466 0.2722
1.5 0.3887 0.3924 0.3175 0.3444
1.7 0.4761 0.4791 0.4200 0.4446
1.9 0.5358 0.5375 0.4883 0.5115
2.0 0.5602 0.5634 0.5148 0.5380
2.5 0.6369 0.6390 0.6027 0.6213
3.0 0.6859 0.6879 0.6566 0.6738
Case 1b Weak correlation ZI 1.1 0.0023 0.0022 0.0022 0.0042
1.2 0.0748 0.0733 0.0622 0.0774
1.3 0.2109 0.2130 0.1531 0.1787
1.4 0.3212 0.3240 0.2415 0.2682
1.5 0.3863 0.3894 0.3138 0.3429
1.7 0.4793 0.4808 0.4187 0.4449
1.9 0.5378 0.5403 0.4886 0.5094
2.0 0.5602 0.5621 0.5118 0.5333
2.5 0.6382 0.6406 0.6011 0.6226
3.0 0.6837 0.6859 0.6556 0.6731
Case 2a Moderate correlation No ZI 1.1 0.0104 0.0102 0.0024 0.0050
1.2 0.3333 0.3429 0.0973 0.1231
1.3 0.4894 0.4973 0.2322 0.2616
1.4 0.5642 0.5684 0.3299 0.3578
1.5 0.6141 0.6186 0.3966 0.4269
1.7 0.6775 0.6796 0.4997 0.5287
1.9 0.7198 0.7212 0.5724 0.5988
2.0 0.7344 0.7351 0.5992 0.6242
2.5 0.7777 0.7789 0.6954 0.7141
3.0 0.8025 0.8034 0.7463 0.7622
Case 2b Moderate correlation ZI 1.1 0.0114 0.0122 0.0021 0.0049
1.2 0.3292 0.3375 0.0969 0.1204
1.3 0.4900 0.4964 0.2345 0.2626
1.4 0.5624 0.5677 0.3265 0.3548
1.5 0.6101 0.6136 0.3938 0.4237
1.7 0.6750 0.6766 0.4975 0.5248
1.9 0.7150 0.7166 0.5681 0.5941
2.0 0.7289 0.7303 0.5980 0.6226
2.5 0.7727 0.7735 0.6911 0.7102
3.0 0.7999 0.8007 0.7447 0.7614
Case 3a Strong correlation No ZI 1.1 0.0114 0.0131 0.0018 0.0039
1.2 0.4044 0.4167 0.0623 0.0852
1.3 0.6199 0.6280 0.1940 0.2290
1.4 0.7176 0.7231 0.3035 0.3437
1.5 0.7660 0.7701 0.3947 0.4325
1.7 0.8212 0.8234 0.5209 0.5537
1.9 0.8515 0.8527 0.5988 0.6270
2.0 0.8629 0.8638 0.6284 0.6571
2.5 0.8929 0.8934 0.7314 0.7531
3.0 0.9072 0.9076 0.7881 0.8055
Case 3b Strong correlation ZI 1.1 0.0145 0.0162 0.0020 0.0040
1.2 0.3987 0.4159 0.0639 0.0881
1.3 0.6192 0.6287 0.1955 0.2312
1.4 0.7147 0.7211 0.3062 0.3455
1.5 0.7658 0.7702 0.3964 0.4342
1.7 0.8231 0.8255 0.5197 0.5520
1.9 0.8537 0.8551 0.5966 0.6261
2.0 0.8633 0.8645 0.6251 0.6529
2.5 0.8931 0.8936 0.7308 0.7536
3.0 0.9076 0.9080 0.7897 0.8071